what is mass of 5 mole of ammonia?calculate the number of NH3 molecules, nitrogen atoms and hydrogen atoms in it



1
Expert's answer
2021-06-21T05:30:18-0400

Molar mass of NH3 = 17.031

5 moles of NH3 = 17.031 × 5

= 85.155g

1 mole = 6.022×1023

Hence, NH3 has 6.022 × 1023 molecules

NH3 has 1N and 3H

Therefore, number of atoms for N = 1/4 × 6.022×1023

= 1.5055×1023atoms

Number of atoms for H = 3/4 × 6.022×1023

= 4.5165 × 1023 atoms